17 September 2014: Wednesday

LDAP:

  • No issues in the update last night.
  • Rolling restart of ldap servers to pick up new set of ACLs.
  • Apparently, yesterday’s update to the daily data deleter code was not quite correct. ┬áRecovered the files AGAIN, and modified the daily data deleter code for hopefully the last time.

Backups:

  • Clones are working their way through, and should finish today or tomorrow.
  • DataDomain continues to be almost full.

Tags:

Leave a Reply

You must be logged in to post a comment.